OpenStudy (juan1857):

Write the equation 7x-2y=2 in polar form

OpenStudy (dumbcow):

x= rcos y = r sin \[7r \cos \theta - 2 r \sin \theta = 2\] solve for r

OpenStudy (juan1857):

I have no idea.

OpenStudy (dumbcow):

try factoring out an "r" from left side

OpenStudy (juan1857):

r(7cos-2sin)=2??